OpenStudy (anonymous):

A man, 2 m tall runs toward a street light at a speed of 3 m/s. If the street light is 10 m tall, how fast is the tip of his shadow moving?

OpenStudy (valpey):

 \[\frac{x}{10} = \frac{3}{8}\] \[x = \frac{30}{8} = 3.75 m/s\]